what is the sum of the interior angles of a polygon with 38 sides.
Answer by dfrazzetto(283)

For any polygon S = (n-2)*180 where n is the number of sides and S is the sum of interior angles.
So:
S = (38-2)*180 = 36*180 = 6480
6480 degrees